If the 4th term of the geometric progression is -32 and the 7th term is 108,what is the 9th term?
Answer by sachi(548)

If the 4th term of the geometric progression is -32 and the 7th term is 108,what is the 9th term?
ans:-
let the 1st term of GP=a & common ratio=r
then 4th term =ar^3= -32 .....eqn 1
and 7th term =ar^6=108,......eqn 2
dividing eqn 2 by eqn 1
we get r^3=108/-32=-27/8......eqn 3
or r=-3/2
now dividing eqn 1 by eqn 3
we will get a= 32*8/27
the 9th term=ar^8= (32*8/27)(-3/2)^8=243 ans